I must decide if the algebraic manipulation is correct or incorrect. if it is incorrect I need to fix it or explaine why it cannot be simplified.
6x-2(x-y)= 6x-2x-2y = 4x-2y

Answer by Alan3354(69443)

I must decide if the algebraic manipulation is correct or incorrect. if it is incorrect I need to fix it or explaine why it cannot be simplified.
6x-2(x-y)= 6x-2x-2y = 4x-2y
--------------
It's incorrect, should be +2y in the center
6x-2(x-y)= 6x-2x+2y = 4x+2y